# A patchy-particle 3-dimensional octagonal quasicrystal

### v1

@@ Abstract
We devise an ideal 3-dimensional octagonal quasicrystal that is based upon the 2-dimensional Ammann-Beenker tiling and that is potentially suitable for realization with patchy particles. Based on an analysis of its local environments we design a binary system of 8- and 5-patch particles that in simulations assembles into a 3-dimensional octagonal quasicrystal. The local structure is subtly different from the original ideal quasicrystal possessing a narrower coordination-number distribution; in fact, the 8-patch particles are not needed and a one-component system of the 5-patch particles assembles into an essentially identical octagonal quasicrystal. We also consider a one-component system of the 8-patch particles; this assembles into a cluster with a number of crystalline domains, but which, because of the coherent boundaries between the crystallites, has approximate eight-fold order. We envisage that these systems could be realized using DNA origami or protein design.
@@
Quasicrystals (QCs) have long range order, as exemplified by the sharp peaks in their diffraction patterns, but without a periodically repeating unit cell. QCs can thus have symmetries that are not possible for periodic crystals. The first example was discovered by Shechtman for a Al/Mn alloy and had icosahedral symmetry [1]. This discovery stimulated a search for further examples and quasicrystalline alloys with decagonal [2], dodecagonal [3, 4] and octagonal [5, 6] symmetry were quickly discovered. More recently, QCs have also been experimentally discovered in soft matter systems [7, 8, 9], but these have been limited to dodecagonal symmetry. Thus, discovering new ways to form QCs and increasing the possible repertoires of structures and symmetries is of significant interest.
@@
Theory and simulations have the potential to significantly contribute to this goal by identifying how interparticle interactions can be designed to achieve quasicrystallinity. QCs involve multiple length scales that are related by a specific irrational ratio, e.g. the golden ratio in the case of icosahedral QCs. This feature is exploited in approaches that use isotropic potentials with multiple length scales [10, 11]. For example, a theoretical framework for designing such ultra-soft potentials has been developed [12, 13, 14]. QCs have also been achieved for multiple length-scale potentials with hard cores albeit more through the systematic exploration of the space of potential parameters, both in 2D [15] and 3D [16, 17].
@@
An alternative approach to design QCs is to attempt to directly design the tendency to form a particular symmetry through directional bonding [18, 19, 20, 21, 22]. The idea is that the directionality helps to induce both local and global non-crystallographic symmetry, where the quasiperiodicity is then a consequence of the formation of that global symmetry. This approach has recently been successfully applied to design patchy particles that form icosahedral quasicrystals in simulations [21, 22]. First, it is important that the patch geometry is designed to be consistent with the desired point group symmetry; e.g. for the icosahedral QC forming systems the patches are directed along subsets of the symmetry axes of $I_{h}$. Although this encourages motifs with the desired local orientational order, it is also important that the system can form a globally ordered structure that is sufficiently stable to outcompete possible periodic crystal forms during assembly. In the icosahedral QC-forming system, this second feature was achieved by basing the patch design on the local environments in an ideal target QC (that was obtained by the cut-and-project approach), thus helping to ensure that the best way that the local environments determined by the patch geometry combine together is as a QC with the target global symmetry.
@@
Octagonal symmetry is one of the more experimentally rarely observed quasicrystalline symmetries with only a few examples in alloys [5, 6]. Moreover, in these examples, the quasicrystalline phase is metastable, being observed to transform to a $\beta$-Mn crystal [23]. The number of examples of octagonal quasicrystal self-assembly in simulations is also modest, all in systems with isotropic interactions, both in 2D [24, 25] and in 3D [17]. Here, the goal is to increase the potential means to realize octagonal QCs by designing a system that can form a 3-dimensional octagonal QC through directional bonding.
@@
Figure 1: (a) The ideal target octagonal quasicrystal. The structure projected down the $c$-axis is that of an Ammann-Beenker tiling with particles at the corners of the square and rhomboidal tiles. The particles are coloured by their coordination number. (b) In 3-dimensions, the square is puckered and the rhombus corresponds to a right- or left-handed helix with a pitch length of the $c$ repeat. (c) The coordination environments in the ideal quasicrystal labelled by their coordination number and their local point group symmetry. The two four-coordinate and six-coordinate environments are enantiomeric. (d) The two patchy particles. (e) The interaction matrix between the four different types of patches on the above two particles.
@@
To apply this approach, we first need an ideal 3D octagonal QC to provide a basis for the patchy-particle designs. Our starting point is the 2D Ammann-Beenker tiling [26] that can, for example, be derived by projection of a 4D hypercubic lattice into 2D (Supplementary Section S2). The tiling consists of squares and rhombi (with internal angles of 45 and 135) with the edges of these polygons being equally likely to be oriented along eight equivalent directions. The simplest way to create a 3D octagonal QC from this tiling would be to place particles at the vertices and have a simple periodic stacking of identical Ammann-Beenker layers (as was done in Ref. [20] for a dodecagonal QC example). However, it is not feasible to realize this structure with patchy particles of well-defined radii and with bonds along the edges of the tiling because the distance across a rhomboidal tile is 0.7654 of the tile edge length and hence would lead to particle overlaps. Instead, we explored displacing particles placed at the vertices of the Ammann-Beenker tiling in $z$ so that only tile edges correspond to bonds. Our solution is presented in Fig. 1. Each bond has a component in the $z$-direction of ±$\pm c/4$ where $c$ is the periodic repeat in that direction. The rhombi become right- or left-handed helices with a pitch length of $c$, whereas the squares remain as a closed circuit of bonds, but where the pairs of diagonally-opposite particles are displaced by $c/4$ (Fig. 1(b)). We choose $c$ so that the distance between particles across the short diagonal of the projected rhombi matches the next-neighbour distance across the diagonals of the squares.
@@
The coordination number distribution for this ideal QC is the same as for the vertices in the Ammann-Beenker tiling with local environments having from 3 to 8 neighbours. Each of the environments is a sub-environment of the 8-coordinate environment, which involves eight-equivalent bonds in a $D_{4d}$ geometry (Fig. 1(c)). The average coordination number is four.
@@
Our basic approach to patchy-particle design is to choose the patch geometry to match the bond directions of an environment [27, 28]. For quasicrystals in particular, we have found that rather than having a different particle type for each environment (e.g. eight in the current case), the complexity of the model can be reduced by representing environments that are subsets of each other by a single patchy particle [21]. A consequence of this approach is that some of the patches would not be involved in bonds in the lower-coordinate environments, and so the particle would have an interaction energy that is lower in magnitude than its possible maximum. A concern might thus be that alternative structures that were better able to utilize all the system’s patches might be more stable.
@@
In the current case, as all the lower-coordinate environments are subsets of the eight-coordinate, one might initially imagine a one-component system of eight-patch particles might be most suitable. However, this would lead to four unused patches per particle in the ideal quasicrystal and so have a significant potential for alternative structure formation. Instead, we initially consider a two-component system that is a mixture of five- and eight-patch particles, where it is envisaged that the eight-patch particles would be used in the 6-, 7- and 8-coordinate environments and the 5-patch particles in the 3-, 4- and 5-coordinate environments. In this case, the composition of 5- to 8-patch particles that would match the ideal quasicrystal is $(7\sqrt{2}-9)/(10-7\sqrt{2}):1\approx 8.95:1$.
@@
Figure 2: (a) Close-up of a cut through a binary octagonal quasicrystal viewed down the 8-fold axis. Bonds are drawn between particles within $5\,\sigma_{\mathrm{LJ}}$ of the cut surface. (b) BOOD for the binary quasicrystal. (c) BOOD for a one-component quasicrystal made of 5-patch particles. BOODs were calculated with the convention that two particles are bonded if the distance between them is lower than 1.5 $\sigma_{\mathrm{LJ}}$ and the interaction energy is lower than -0.2 $\epsilon_{\mathrm{LJ}}$. BOODs are projected onto a plane perpendicular to the 8-fold symmetry axis using Lambert projection. (d) Diffraction patterns of the binary octagonal QC viewed down the 8-fold axis and a 2-fold axis. (e) Radial distribution functions for the assembled binary, P5, and P8 clusters and the ideal OQC and $C2/c$ crystal. (f) Coordination number distributions for the same systems as (e) (calculated using the same energy criterion as the BOODs in the assembled QCs).
@@
All the patches of the 8-patch particle are equivalent by symmetry, whereas the 5-patch particle has three sets of non-equivalent patches. The specificity of the patch-patch interactions can be used to help favour the formation of the target structure. In this case, we only allow patches to interact if they would be involved in bonds in the ideal octagonal QC (the resulting interaction matrix is shown in Fig. 1(e)). For example, there are never any bonds between 6-, 7- and 8-coordinate environments in the ideal octagonal QC, so the patches on the 8-patch particle are non self-interacting.
@@
To simulate the patchy particles, we use a modified Lennard-Jones potential where the attractive component is modulated by angular and torsional factors such that the full attractive interaction is only obtained if interacting patches point directly at each other and the particles have the correct relative orientation [27, 29, 28]. The simulations were performed with a GPU-enabled Monte Carlo algorithm [30]. The full form of the potential, the particle designs and further details about the simulations are given in the Supplementary Information.
Fig. 2(a) shows a cut through a large binary cluster grown in our simulations. From the structure it can be clearly seen that bonds are oriented along eight equivalent directions. This is confirmed by the bond-orientational order diagram (BOOD) (Fig. 2(b)) which has clear eight-fold symmetry. Similarly, the diffraction pattern (Fig. 2(d)) also exhibits eight-fold symmetry and quasiperiodic character, e.g. the ratio of the positions of the first two peaks in the $x$ direction is $\sqrt{2}$. The cluster is an octagonal quasicrystal.
In Fig. 2(f) we show the coordination number distribution for the assembled cluster. The average coordination number of our assembled octagonal QC is close to four ($\langle CN\rangle$ =3.81; see Table S6), i.e. nearly the same as the ideal QC. However, there is a relative lack of high-coordinate particles compared to the ideal Ammann-Beenker structure. Only about 1% of particles have a coordination number of six or more even though the fraction of 8-patch particles is 9%. This feature suggests that the 8-patch particles may not be actually needed for quasicrystal formation. We therefore ran simulations of a one-component system of 5-patch particles (with the same patch-patch interaction matrix (Fig. 1(e))). An octagonal quasicrystal again formed (Figs. S2, S3)) that had effectively identical structural properties to the binary system, e.g. the coordination number distribution (Fig. 2(f)) and radial distribution functions are very similar (Fig. 2(e)).

### v2

@@ Abstract
We devise an ideal 3-dimensional octagonal quasicrystal that is based upon the 2-dimensional Ammann-Beenker tiling and that is potentially suitable for realization with patchy particles. Based on an analysis of its local environments we design a binary system of 5- and 8-patch particles that in simulations assembles into a 3-dimensional octagonal quasicrystal. The local structure is subtly different from the original ideal quasicrystal possessing a narrower coordination-number distribution; in fact, the 8-patch particles are not needed and a one-component system of the 5-patch particles assembles into an essentially identical octagonal quasicrystal. We also consider a one-component system of the 8-patch particles; this assembles into a cluster with a number of crystalline domains, but which, because of the coherent boundaries between the crystallites, has approximate eight-fold order. We envisage that these systems could be realized using DNA origami or protein design.
I Introduction
@@
Quasicrystals (QCs) have long-range order, as exemplified by the sharp peaks in their diffraction patterns, but without a periodically repeating unit cell. QCs can thus have symmetries that are not possible for periodic crystals. The first example was discovered by Shechtman for an Al/Mn alloy and had icosahedral symmetry.[1] This discovery stimulated a search for further examples and quasicrystalline alloys with decagonal,[2] dodecagonal[3, 4] and octagonal[5, 6] symmetry were quickly discovered. More recently, QCs have also been experimentally discovered in soft matter systems,[7, 8, 9] but these have been generally limited to dodecagonal symmetry. Thus, discovering new ways to form QCs and increasing the possible repertoires of structures and symmetries is of significant interest.
@@
Theory and simulations have the potential to significantly contribute to this goal by identifying how interparticle interactions can be designed to achieve quasicrystallinity. QCs involve multiple length scales that are related by a specific irrational ratio, e.g. the golden ratio in the case of icosahedral QCs. This feature is exploited in approaches that use isotropic potentials with multiple length scales.[10, 11] For example, a theoretical framework for designing such ultra-soft potentials has been developed.[12, 13, 14] QCs have also been achieved for multiple length-scale potentials with hard cores, albeit more through the systematic exploration of the space of potential parameters, both in 2D[15] and 3D.[16, 17, 18] Inverse design methods also hold significant promise for quasicrystal discovery.[19]
@@
An alternative approach to design QCs is to attempt to directly design the tendency to form a particular symmetry through directional bonding.[20, 21, 22, 23, 24, 25] The idea is that the directionality helps to induce both local and global non-crystallographic symmetry, where the quasiperiodicity is then a consequence of the formation of that global symmetry. This approach has recently been successfully applied to design patchy particles that form icosahedral quasicrystals in simulations.[23, 24] First, it is important that the patch geometry is designed to be consistent with the desired point group symmetry; e.g. for the icosahedral QC-forming systems the patches are directed along subsets of the symmetry axes of $I_{h}$. Although this encourages motifs with the desired local orientational order, it is also important that the system can form a globally ordered structure that is sufficiently stable to out-compete possible periodic crystal forms during assembly. In the icosahedral QC-forming system, this second feature was achieved by basing the patch design on the local environments in an ideal target QC (that was obtained by the cut-and-project approach), thus helping to ensure that the best way that the local environments determined by the patch geometry combine together is as a QC with the target global symmetry.
@@
Octagonal symmetry is one of the more experimentally rarely observed quasicrystalline symmetries with only a few examples in alloys.[5, 6] Moreover, in these examples, the quasicrystalline phase is metastable, being observed to transform to a $\beta$-Mn-type crystal.[26] The number of examples of octagonal quasicrystal self-assembly in simulations is also modest, all in systems with isotropic interactions, both in 2D[27, 28] and in 3D.[17] Here, the goal is to increase the potential means to realize octagonal QCs by designing a system that can form a 3-dimensional octagonal QC through directional bonding.
@@
Figure 1: The ideal target octagonal quasicrystal. (a) The structure projected down the $c$-axis is that of an Ammann-Beenker tiling with particles at the corners of the square and rhomboidal tiles. The particles are coloured by their coordination number (See Fig. 2). (b) A side-view along one of the two-fold axes. (c) In 3-dimensions, the square is puckered and the rhombus corresponds to a right- (illustrated) or left-handed helix with a pitch length of the $c$ repeat.
Figure 2: (a) The coordination environments in the ideal quasicrystal labelled by their coordination number and their local point group symmetry. The two four-coordinate and six-coordinate environments are enantiomeric. (b) The two patchy particles. (c) The interaction matrix between the four different types of patches on the above two particles.
II An ideal octagonal quasicrystal
@@
To apply this approach, we first need an ideal 3D octagonal QC to provide a basis for the patchy-particle designs. Our starting point is the 2D Ammann-Beenker tiling [29] that can, for example, be derived by projection of a 4D hypercubic lattice into 2D (Supplementary Section S1). The tiling consists of squares and rhombi (with internal angles of 45 and 135) with the edges of these polygons being equally likely to be oriented along eight equivalent directions. The simplest way to create a 3D octagonal QC from this tiling would be to place particles at the vertices and have a simple periodic stacking of identical Ammann-Beenker layers (as was done in Ref. 22 for a dodecagonal QC example). However, it is not feasible to realize this structure with patchy particles of well-defined radii and with bonds along the edges of the tiling because the distance across the short diagonal of a rhomboidal tile is 0.7654 of the tile edge length and hence would lead to particle overlaps. Instead, we explored how to place particles at the vertices of the Ammann-Beenker tiling at different heights in $z$ so that only tile edges correspond to bonds and there are no particle overlaps. Our solution is presented in Fig. 1. Each bond has a component in the $z$-direction of $\pm c/4$ where $c$ is the periodic repeat in that direction. The rhombi become right- or left-handed helices with a pitch length of $c$, whereas the squares remain as a closed circuit of bonds, but where the pairs of diagonally-opposite particles are displaced by $c/4$ (Fig. 1(c)). We choose $c$ so that the distance between particles across the short diagonal of the projected rhombi matches the next-neighbour distance across the diagonals of the squares.
@@
The coordination number distribution for this ideal QC is the same as for the vertices in the Ammann-Beenker tiling with local environments having from 3 to 8 neighbours. Each of the environments is a sub-environment of the 8-coordinate environment, which involves eight-equivalent bonds in a $D_{4d}$ geometry (Fig. 2(a)). The average coordination number is four.
The above proposed decoration of the square and two rhomboidal cells is clearly compatible with quasicrystal formation in systems made up of such cells, but it is interesting to ask how such a decoration constrains the ways that these cells could potentially combine. The decoration induces an edge-matching rule between the three types of cells (Fig. S2). Although this differs from the edge-matching rule of the Ammann-Beenker tiling that when combined with the appropriate vertex-matching rule ensures quasiperiodicity,[30] this is not necessarily a problem for our approach, because, like in previous patchy-particle quasicrystals,[23, 24] entropy maximization is expected to be the driving force for any quasiperiodicity observed.
This ideal QC is significantly different from previously observed 3-dimensional octagonal QCs,[5, 6, 17] as although they can also be analysed in terms of square-rhomb tilings, the decoration of the tiles with particles is much more complex, with most models of the experimental examples being related to the $\beta$-manganese structure.[31, 32] However, our ideal octagonal QC does have some structural similarities to the 3-dimensional decagonal QC observed in simulations in Ref. 17 for a multi-minimum potential in that this quasicrystal is also based on a classical quasiperiodic tiling, namely the Tübingen tiling, with particles at the vertices of the tiles and one of the tiles (the pentagon) being associated with helical configurations.
III Patchy-particle design
@@
Our basic approach to patchy-particle design is to choose the patch geometry to match the bond directions of an environment. [33, 34] For quasicrystals in particular, we have found that rather than having a different particle type for each environment (e.g. eight in the current case), the complexity of the model can be reduced by representing environments that are subsets of each other by a single patchy particle.[23, 24] A consequence of this approach is that some of the patches would not be involved in bonds in the lower-coordinate environments, and so the particle would have an interaction energy that is lower in magnitude than its possible maximum. A concern might thus be that alternative structures that were better able to utilize all the system’s patches might be more stable.
@@
In the current case, as all the lower-coordinate environments are subsets of the eight-coordinate environment, one might initially imagine that a one-component system of eight-patch particles might be most suitable. However, this would lead to an average of four unused patches per particle in the ideal quasicrystal and so have a significant potential for alternative structure formation. Instead, we initially consider a two-component system that is a mixture of five- (P5) and eight-patch (P8) particles, where it is envisaged that the eight-patch particles would be used in the 6-, 7- and 8-coordinate environments and the 5-patch particles in the 3-, 4- and 5-coordinate environments. In this case, the composition of 5- to 8-patch particles that would match the ideal quasicrystal is $(7\sqrt{2}-9)/(10-7\sqrt{2}):1\approx 8.95:1$ (Table S2).
@@
Figure 3: (a) Close-up of a cut through a binary octagonal quasicrystal viewed down the 8-fold axis. Bonds are drawn between particles within $5\,\sigma_{\mathrm{LJ}}$ of the cut surface, and these particles are visualized as small green (P5) or cyan (P8) spheres. Particles further away from this cut plane are visualized as larger grey spheres. The yellow ellipse highlights an “overlapping-squares” motif (see Fig. 4(a)). (b) BOOD for the binary quasicrystal. (c) BOOD for a one-component quasicrystal made of 5-patch particles. (d) Diffraction patterns of the binary octagonal QC viewed down the 8-fold axis and a 2-fold axis. (e) Radial distribution functions for the assembled binary, P5, and P8 clusters and the ideal OQC and $C2/c$ crystal. (f) Coordination number distributions for the same systems as (e) (calculated for the assembled QCs using the same energy and distance criteria as for the BOODs).
@@
All the patches of the 8-patch particle are equivalent by symmetry, whereas the 5-patch particle has three sets of non-equivalent patches. The specificity of the patch-patch interactions can be used to help favour the formation of the target structure. In this case, we only allow patches to interact if they would be involved in bonds in the ideal octagonal QC (the resulting interaction matrix is shown in Fig. 1(e)). For example, there are never any bonds between 6-, 7- and 8-coordinate environments in the ideal octagonal QC, so the patches on the 8-patch particle are not self-interacting. Similarly, patch 2 on the 5-patch particle would never bond to itself in the ideal quasicrystal, so we also make this patch non self-interacting.
IV Methods
IV.1 Patchy-particle potential
@@
To simulate the patchy particles, we use a modified Lennard-Jones potential where the attractive component is modulated by angular and torsional factors such that the full attractive interaction is only obtained if interacting patches point directly at each other and the particles have the correct relative orientation.[33, 35, 34] The specific form of the interactions between the patchy particles is the same as in Refs. 35, 34. The interaction is based on a cut-and-shifted Lennard-Jones potential $V_{\mathrm{LJ}}^{\prime}$ where
$V_{\mathrm{LJ}}^{\prime}(r)=\begin{cases}V_{\mathrm{LJ}}(r)-V_{\mathrm{LJ}}(r_{\mathrm{cut}})&:r<r_{\mathrm{cut}}\\ 0&:r\geq r_{\mathrm{cut}}\end{cases}$ (1)
and
$V_{\mathrm{LJ}}(r)=4\varepsilon_{\mathrm{LJ}}\left[\left(\frac{\sigma_{\mathrm{LJ}}}{r}\right)^{12}-\left(\frac{\sigma_{\mathrm{LJ}}}{r}\right)^{6}\right].$ (2)
The patchy particle potential is described by a Lennard-Jones repulsive core and an attractive tail that is modulated by angular and torsional dependent functions:
$V_{ij}(\mathbf{r}_{ij},\mathbf{\Omega}_{i},\mathbf{\Omega}_{j})=\begin{cases}V_{\mathrm{LJ}}^{\prime}(r_{ij})&:r_{ij}<\sigma_{\mathrm{LJ}}^{\prime}\\ V_{\mathrm{LJ}}^{\prime}(r_{ij})\underset{{\mathrm{patch~pairs~}\alpha,\beta}}{\max}\left[\varepsilon_{\alpha\beta}V_{\mathrm{ang}}(\mathbf{\hat{r}}_{ij},\mathbf{\Omega}_{i},\mathbf{\Omega}_{j})V_{\mathrm{tor}}(\mathbf{\hat{r}}_{ij},\mathbf{\Omega}_{i},\mathbf{\Omega}_{j})\right]&:r_{ij}\geq\sigma_{\mathrm{LJ}}^{\prime}\end{cases},$ (3)
where $\mathbf{r}_{ij}$ is the interparticle vector, $\alpha$ and $\beta$ are patches on particles $i$ and $j$ respectively, $\mathbf{\Omega}_{i}$ is the orientation of particle $i$, $\sigma_{\mathrm{LJ}}^{\prime}$ corresponds to the distance at which $V_{\mathrm{LJ}}^{\prime}$ passes through zero and $\varepsilon_{\alpha\beta}$ is a measure of the relative strength of the interactions between patches $\alpha$ and $\beta$. We set the cutoff distance $r_{\mathrm{cut}}=2.5\,\sigma_{\mathrm{LJ}}$.
The angular modulation term $V_{\mathrm{ang}}$ is a measure of how directly the patches $\alpha$ and $\beta$ point at each other, and is given by
$V_{\mathrm{ang}}(\mathbf{\hat{r}}_{ij},\mathbf{\Omega}_{i},\mathbf{\Omega}_{j})=\exp\left(-\frac{\theta_{\alpha ij}^{2}}{2\sigma_{\mathrm{ang}}^{2}}\right)\exp\left(-\frac{\theta_{\beta ji}^{2}}{2\sigma_{\mathrm{ang}}^{2}}\right).$ (4)
$\theta_{\alpha ij}$ is the angle between the patch vector $\mathbf{\hat{P}}_{i}^{\alpha}$, representing the patch $\alpha$ on particle $i$, and $\mathbf{\hat{r}}_{ij}$. $\sigma_{\mathrm{ang}}$ is a measure of the angular width of the patch.
The torsional modulation term $V_{\mathrm{tor}}$ describes the variation in the potential as either of the particles is rotated about the interparticle vector $\mathbf{r}_{ij}$ and is given by
$V_{\mathrm{tor}}(\mathbf{\hat{r}}_{ij},\mathbf{\Omega}_{i},\mathbf{\Omega}_{j})=\exp\left(-\frac{1}{2\sigma_{\mathrm{tor}}^{2}}\left[\min\limits_{\phi^{\mathrm{offset}}_{\alpha\beta}}\left(\phi_{\alpha\beta}-\phi^{\mathrm{offset}}_{\alpha\beta}\right)\right]^{2}\right).$ (5)
where $\phi^{\mathrm{offset}}$ is the preferred value of the torsional angle $\phi$. To define the torsional angle $\phi_{\alpha\beta}$, a unique reference vector is associated with each patch. In order to capture the symmetry of an environment, more than one equivalent offset angle can be defined, in which case we find the minimum value of $\phi-\phi_{\mathrm{offset}}$ across the set of equivalent offset angles. The torsional interaction prevents free rotation about a patchy bond and encourages particles to bind with the correct relative orientation, thus facilitating the propagation of long-range orientational order.
As in previous work we choose $\sigma_{\mathrm{tor}}=2\,\sigma_{\mathrm{ang}}$. We also choose $\sigma_{\mathrm{ang}}$ to be sufficiently narrow to give strong directional bonding, but not too narrow as to hinder the kinetics. We use both $\sigma_{\mathrm{ang}}=0.2$ and 0.3 radians. $\varepsilon_{\alpha\beta}$ is either 0 or 1, matching the interaction matrix in Fig. 1(e); as this simple choice was successful in achieving quasicrystal assembly, there was no attempt to further optimize the strengths of the interacting patches. We use $\sigma_{\mathrm{LJ}}$ (the distance at which the Lennard–Jones potential is zero) as our unit of length, and the Lennard–Jones well depth $\varepsilon_{\mathrm{LJ}}$ as our unit of energy. Temperatures are given in reduced form, i.e. $T^{*}=k_{B}T/\varepsilon_{\mathrm{LJ}}$.
The details of the patchy particle designs for the binary system proposed in Section III are given in Table S4. The symmetry axis of the 8-patch particle is used as the reference vector for the torsional component of the potential. An offset angle of 180 ensures that the preferred relative geometry is with the eight-fold axes of the particles alligned. The patches of the 5-patch particle are a subset of those of the 8-patch particle.
IV.2 Simulations
The simulations were performed with a GPU-enabled Monte Carlo algorithm.[36] Our basic simulation protocol is similar to our previous work on 3-dimensional patchy-particle quasicrystals.[23, 22, 24] Namely, assembly is started from a low-density fluid (with 20 000 particles), choosing a temperature that is just sufficiently low enough to allow nucleation of the ordered phase (we generally want to avoid multiple nucleation events). The composition of the fluid in the binary system was chosen to match that for the ideal target quasicrystal. Once we have grown a moderately-sized cluster (typically of order 10 000 particles) we then place this cluster in a larger box (and with a larger reservoir of particles in the fluid phase) and continue growth. We may increase the temperature slightly in this second phase to avoid any further nucleation events. For the results reported in Section V.1 and V.2, we typically ran the simulations until a cluster with close to 100 000 particles was obtained. More details of these simulations are given in Table 1.
For the simulations reported in Section V.3, which test the effects of varying potential and design parameters, instead the clusters were usually grown just until they were sufficiently large for QC formation to be confidently determined from the diffraction patterns and BOODs; this corresponded to at least 10 000 particles.